OpenStudy (sleepyjess):

Use basic identities to simplify the expression. \(\dfrac{cos^2\theta}{sin^2\theta}~+~csc\theta~sin\theta\) \(csc^2\theta\) \(sec^2\theta\) 1 \(tan^2\theta\)

OpenStudy (sleepyjess):

@satellite73 @e.mccormick

OpenStudy (anonymous):

\[ \dfrac{cos^2\theta}{sin^2\theta}~+~csc\theta~sin\theta = \dfrac{cos^2\theta}{sin^2\theta}~+~ \frac{sin(\theta)}{sin(\theta)} = cot^2(\theta)+1 = csc(\theta) \]

OpenStudy (sleepyjess):

But that isn't one of the options?

OpenStudy (sleepyjess):

@myininaya

OpenStudy (e.mccormick):

Um, is that supposed to be csc on the right?

OpenStudy (anonymous):

Sorry, I missed the squared off there, it should be \[csc^2(\theta)\]

OpenStudy (e.mccormick):

Yah.

OpenStudy (e.mccormick):

Well, there is more than one possible answer. They may have it in a different form.